# PromptRPA: Generating Robotic Process Automation on Smartphones from Textual Prompts

### v1

@@ 6. User Evaluation
Given the influence of user feedback on system performance, we assessed PromptRPA’s usability and robustness through engagement with diverse users.
@@ 6.1. Experiment Setup
Participant Demographics: We recruited a total of 12 participants for our study. Their proficiency was determined through a pre-study questionnaire that evaluated their comfort level with smartphone usage, experience with RPA tools, and frequency of smartphone use. The participants were divided into the following two categories according to their proficiency.
@@
  • Unskilled Users: 6 participants were less proficient in smartphone usage (4 females; aged from 50 to 69, M = 60.00). The group included elders, beginners, or infrequent smartphone users.
@@
  • Predefined Tasks: The 4 participants within a set are assigned the same 8 predefined tasks, split across 2 applications: one commonly used and one less commonly used. For each application, 2 common and 2 uncommon RPA tasks are chosen, based on criteria detailed in Section 5.1. Part of the tasks are shown in Table 2. In each four-participant set, we assume that ‘a’ and ‘b’ are skilled users, whereas ‘c’ and ‘d’ are unskilled users. To ensure balance, for each application, one common task is performed manually by ‘a’ and ‘c’, while ‘b’ and ‘d’ use PromptRPA. For the other common task, roles are reversed: ‘a’ and ‘c’ use PromptRPA, while ‘b’ and ‘d’ perform it manually. The approach is similar for uncommon tasks.
  • User-Suggested Tasks: In addition to the predefined tasks, each participant is allowed to propose 2 RPA tasks for each of their assigned applications. We encouraged participants to suggest challenging tasks. For these 2 tasks, participants will employ both manual execution and PromptRPA. To counterbalance the effects of the method sequence, the order of methods is alternated across tasks.
Methods: Almost all tasks in the experiment, requiring in-depth application functionality, have been validated as unsupported by Google Assistant and HUAWEI Celia, which excel in information gathering but fail to execute tasks. For manual execution, participants are allowed to seek assistance from online resources and people around them. For tasks carried out using PromptRPA, participants can freely provide any information through textual prompts, which can either describe the task’s goal or outline the detailed steps required for the RPA task. After completing the RPA task, participants are required to use a different expression to invoke the RPA task again.
Metrics: We recorded both the time and success rate for each task, irrespective of whether it was performed manually or via PromptRPA. For tasks carried out using PromptRPA, the frequency of user interventions was also recorded.
Experimental Device: Participants used their own Android smartphones for the experiment. PromptRPA’s client is installed on these devices.
In summary, we collected 192 ($12\times 8\times 2$) data points for tasks completed using PromptRPA (half of which represent first-time usage and the other half for second-time usage), and 96 ($12\times 8$) data points for tasks completed manually.
@@ 6.2. Experimental Results
This section elucidates the experiment’s results, focusing on metrics including success rate, completion time, and frequency of user interventions, as well as subject feedback.
Table 8. Impact of different methods and participant groups on task success rate (%).
Manual PromptRPA(First Time) PromptRPA(Second Time)
@@
Success Rate. Table 8 reveals that PromptRPA outperforms manual methods in both skilled and unskilled groups in terms of task success rate. Failures during the first use of PromptRPA were largely attributable to the low quality of online tutorials for uncommonly used applications. Failures during the second attempt were mainly due to page data loading issues that blocked access to operation widgets.
@@
Completion Time. Figure 14 shows that completing tasks with PromptRPA takes significantly longer during the first use than when carried out manually (p-value<0.001). The time stems primarily from the system’s retrieval for online tutorials and exploration of on-screen operation sequences. Figure 14 indicates that the completion time correlates positively with the number of steps involved during the first use of PromptRPA, which is more pronounced than with manual methods and the second usage of PromptRPA. This latency is tied to the efficiency limitations of the LLM. Notably, time decreases substantially on second use, as the system’s historical RPA repository eliminates the need for additional tutorial retrieval and exploration, thereby becoming significantly faster than manual methods (p-value<0.001).
@@
User Intervention. When using PromptRPA, each new task required 1.92 user interventions (sd = 1.07) on average. Of these, 0.75 interventions per task (sd = 1.04) occurred during the operation mapping. The Assessment Agent accurately predicted the need for these interventions with a precision rate of 66.67% and a recall rate of 83.33%. Notably, the second time a task was executed using PromptRPA, user interventions were no longer required.
@@
Figure 15. Distribution of participant ratings for each statement in Table 10, assessing various aspects of the system. Ratings range from 1 (strongly disagree) to 7 (strongly agree).
Subjective Feedback. Post-experiment interviews were conducted with participants, and their 7-point Likert scale responses are summarized in Figure 15.
  • Low Effort: Participants generally found the system’s required effort reasonable (5.42 in 7). They commonly noted that most interventions were for locating small icons. P3 expressed surprise and satisfaction, “Once I taught it the process, next time I only needed to provide a few keywords to complete the task.” However, P2 expressed a desire for more automation to reduce mental effort.
  • Ease of Use: With a high average score of 6.33, participants found PromptRPA user-friendly, helping them complete tasks they otherwise could not. P1 suggested incorporating images into the step descriptions for clarity, similar to the online tutorials including screenshots. P4, an older user, remarked, “Adding steps is quite challenging for me; I hope the tutorials could be more accurate.”
  • Low Time Pressure: Participants generally understood the longer initial setup time but expressed hopes for future improvements (5.83 in 7). P5 noted, “Although the process took a bit of time, the real-time feedback made it feel like a learning experience, which was acceptably paced for me.”
@@
  • Decision and Interactive Intelligence: Users awarded high scores of 6.50 and 6.08, respectively, commending these aspects of the system. P3 specifically praised the feature that “aligns the smartphone’s operational page with each step description”. Similarly, P5 stated, “Before, I struggled with the tutorial’s first step not matching my current page. Now it is a significant improvement.”. However, room for improvement was noted. P9 advocated for a more “aggressive” execution approach, arguing the current system is slightly conservative. P11 noted that some steps for confirmation were unnecessary. P12 observed, “Tasks generally succeeded, but some took a roundabout path.”
  • Interactive Flexibility: Participants felt unrestricted in expressing queries (6.67 in 7). P1 found the context supplement very helpful, which allows for queries without the need for “careful consideration”. Besides textual expression, P3 expressed a desire for more intuitive communication methods, “Guess what I am trying to do and help me out when I am stuck in a certain step.”
@@
  • Satisfaction and Willingness to Reuse: Overall, participants were very satisfied with PromptRPA (6.25 in 7) and willing to reuse (6.58 in 7). Many participants praised the system’s transparency and user-centered approach. P8 stated, “It gives me a feeling like I’m navigating through the tasks. It’s very interesting.” The utility of PromptRPA in simplifying otherwise complex tasks was another strong point. P1 expressed, “I often ask my family how to use various features on my phone, but it’s challenging to follow their instructions. Now I can just copy their responses and the system executes the steps for me.” Similarly, P12 shared, “I find new applications, especially those with complex functions, very daunting. It’s such a hassle to contact customer service and then follow their steps. PromptRPA saves me from all that trouble.” P9 pointed out, “I can define a few routine tasks, like accounting after spending money. The agent can then automate these tasks for me.” P3 stated, “I often forget the steps, so the replay is very useful for me.”
@@ 7. Discussion
In this section, we delve deeper into the significance and applications of our proposed PromptRPA. We also discuss the limitations of the system and potential future work.
7.1. Physical Path-finding v.s. Digital Navigation
We draw an analogy between physical path-finding and digital navigation to elucidate PromptRPA’s implications. This comparison helps highlight the universal principles embedded in PromptRPA, applicable in both physical and digital realms.
Mapping Out Task Knowledge: Just as a physical journey begins with map consultation, PromptRPA starts its digital journey with information collection. By merging existing and new knowledge, it provides a detailed task overview, similar to how a map provides guidance.
Plotting the Expedition: A traveler uses a map to chart a route, marking key waypoints. Similarly, during instruction generation, our system transforms complex data into concise, executable instructions.
Embarking on the Journey: Physically traveling a planned route parallels our system’s operation mapping, where plans are not only devised but actualized, dynamically adapting to the digital landscape’s ever-changing terrain.
In essence, PromptRPA introduces a general digital navigation framework. It aims to convert information across modalities (from texts to operations), with its stages focusing on bridging gaps between them and minimizing information loss. Therefore, each stage can be further refined with domain-specific enhancements, with its multi-agent framework providing great convenience.
@@ 7.2. Open and Accumulated Knowledge
In addition to the collaboration of its multi-agents, PromptRPA’s effectiveness is derived from its open and accumulated knowledge, ensuring its broad applicability and efficiency.
PromptRPA leverages massive online tutorials to address a broader range of user needs, surpassing existing work that relies on limited datasets (Li et al., 2020a; Venkatesh et al., 2022), and eliminating the need for prior data collection (Vu et al., 2023). It supports functionalities deep within applications, which exceeds the capabilities of current assistants like Google Assistant and HUAWEI Celia. Moreover, PromptRPA’s approach to mobile semantics is open-ended, enabling exploration beyond rigid adherence to given instructions. The Mobile Semantics Agent interprets real-time page contexts, and in conjunction with existing instructions, the Grounding Agent predicts the operation widgets. The open-ended approach allows PromptRPA to handle diverse textual inputs and adapt to changing environments, significantly expanding RPA’s application scope.
PromptRPA’s knowledge base evolves through user usage and feedback. The evolution manifests in improved success rates, rising from 72.80% to 95.21%, and reduced completion time for repeated tasks (Figure 14). Notably, the accumulation of knowledge contributes to a reduction in user interventions. When similar task requirements arise again, PromptRPA operates autonomously, only requiring user intervention for contextual parameter adjustments or significant interface updates. The accumulated knowledge empowers PromptRPA to tackle increasingly complex and nuanced tasks, dynamically adapting to emerging user needs.
@@ 7.3. Applications
By incorporating a vast array of external knowledge and accepting various textual inputs, PromptRPA supports a wide range of task executions, benefiting both end users and developers.
For end users, PromptRPA improves the usability of smartphones and facilitates assistance between peers. It offers guided support for users navigating unfamiliar smartphone functions, enabling them to access and automate tasks deep within applications. Its capacity to handle a broader range of tasks allows users to personalize their smartphone experience by automating their own complex or repetitive tasks. Furthermore, textual assistance from various sources, such as family, friends, or online customer service, can be fed to PromptRPA for automating tasks, which is valuable for teamwork or external support. For instance, an elderly user receiving textual assistance from family members can simply input the text into PromptRPA for automating tasks, eliminating the need for constant reference to instructions and navigating through complex interfaces.
For developers, PromptRPA can reduce the costs of developing and maintaining task automation. It can cost-effectively transform a vast array of existing tutorial texts into operations, seamlessly integrated into current voice assistants or online customer service, thereby expanding the range of supported tasks. Application and platform developers can leverage PromptRPA to transform text-based cases in bulk into interactive tutorials or executable test scripts. It significantly reduces the cost of updating voice assistants, interactive tutorials, and test scripts, as developers need only update the source texts regularly to reflect the latest application functions and interfaces. Nonetheless, in sensitive applications such as banking, implementing robust authentication mechanisms is crucial to prevent malicious command execution.
@@ 7.4. Limitations and Future Work
We outline the limitations of PromptRPA, along with possible directions for future enhancements.
Adapting Dialogue Mechanisms for Diverse User Groups: The current design of the Analysis Agent offers a generalized approach without catering to the specific needs of different user demographics. For instance, older and younger users may articulate issues differently, necessitating a more nuanced, tailored dialogue system to accommodate their needs.
Parsing Complex Tutorial Logic: While our system can parse nested tutorial texts, there is room for improvement. For example, automating responses to conditional branches like “If the version number is less than 14.0, please upgrade the app” could enhance the system’s efficiency. Future work will focus on more sophisticated logic analysis, including nested conditions.
Improving Semantic Understanding of User Interfaces: In the present design, the Mobile Semantics Agent emphasizes the textual descriptions of mobile pages. However, current models often fall short, particularly when faced with pages not meeting accessibility standards. It results in a notable loss of information during the translation from interface visuals to text. Moving forward, we aim to enhance algorithms that better recognize semantic structures within GUI, with the potential improvement of image recognition technologies.
Balancing Comprehensive Coverage with In-Depth Improvement: Currently, in order to expand the application scope, PromptRPA applies the capabilities of LLM across various stages. However, it has not deeply explored some domains, such as addressing ambiguities in natural language and setting instruction parsing templates to reduce the risk of LLM fabrication. As PromptRPA’s usage data increases, we will systematically collect domain-specific data and user feedback to fine-tune the LLM or train local models, gradually reducing reliance on generalized templates and increasing its capacity for nuanced, context-aware processing.

### v2

@@ 6. User Evaluation
Given the influence of user intervention on system performance, we assessed Prompt2Task’s usability through engagement with diverse users.
@@ 6.1. Experiment Setup
Participant Demographics: We recruited a total of 12 participants for our study through online postings in local communities. Each participant was compensated at a rate of \$15 per hour for their contribution. Before the study began, each participant was required to complete an informed consent form that detailed the nature of the study, the extent of data collection, and the installation of the Prompt2Task client on their personal devices (detailed in Section 4.3). We confirmed that no personal data would be retained post-study without explicit permission. This study has been reviewed and approved by the appropriate Institutional Review Board (IRB).
The participant’s proficiency was determined through a pre-study questionnaire that evaluated their comfort level and frequency with smartphone usage. The participants were divided into the following two categories according to their proficiency.
@@
  • Unskilled Users: 6 participants were less proficient in smartphone usage (4 females; aged from 50 to 69, M = 60.00). They are elders, beginners, or infrequent smartphone users.
@@
  • Predefined Tasks: The 4 participants within a set are assigned the same 8 predefined tasks, split across 2 applications: one commonly used and one less commonly used. For each application, 2 common and 2 uncommon automation tasks are chosen, based on criteria detailed in Section 5.1. Part of the tasks are shown in Table 3. In each four-participant set, we assume that ‘a’ and ‘b’ are skilled users, whereas ‘c’ and ‘d’ are unskilled users. To ensure balance, for each application, one common task is performed manually by ‘a’ and ‘c’, while ‘b’ and ‘d’ use Prompt2Task. For the other common task, roles are reversed: ‘a’ and ‘c’ use Prompt2Task, while ‘b’ and ‘d’ perform it manually. The approach is similar for uncommon tasks.
  • User-Suggested Tasks: In addition to the predefined tasks, each participant is allowed to propose 2 automation tasks for each of their assigned applications. We encouraged participants to suggest challenging tasks. For these tasks, participants will employ both manual execution and Prompt2Task. To counterbalance the effects of the method sequence, the order of methods is alternated across tasks.
Methods: Almost all tasks in the experiment, requiring in-depth application functionality, have been validated as unsupported by Google Assistant and HUAWEI Celia, which excel in information gathering but fail to execute tasks. For manual execution, participants are allowed to seek assistance from online resources and people around them. For tasks carried out using Prompt2Task, participants can freely provide any information through textual prompts, which can either describe the task’s goal or outline the detailed steps required for the task. After completing the task, participants are required to use a different expression to invoke the task again.
Metrics: We recorded the time and success rate for each task, both when performed manually and via Prompt2Task. For tasks carried out using Prompt2Task, the frequency of user interventions was also recorded.
Experimental Device: Participants used their own Android smartphones for the experiment. Prompt2Task’s client is installed on these devices (detailed in Section 4.3).
In summary, we collected 192 ($12\times 8\times 2$) data points for tasks completed using Prompt2Task (half of which from first-time usage and the other half from second-time usage), and 96 ($12\times 8$) data points for tasks completed manually.
@@ 6.2. Experimental Results
This section demonstrates the experiment’s results, focusing on metrics including success rate, completion time, and frequency of user interventions, as well as subjective feedback.
Table 12. The task success rate (%) of different methods and participant groups. Tasks are considered successful only if completed within 10 minutes. “Prompt2Task (Second Time)” refers to tasks that have been previously executed and are being re-executed with a different textual prompt.
Manual Prompt2Task (First Time) Prompt2Task (Second Time)
@@
Success Rate. Table 12 reveals that Prompt2Task outperforms manual methods in both skilled and unskilled groups in terms of task success rate. Failures during the first use of Prompt2Task were largely attributable to the low quality of online tutorials for uncommonly used applications. Failures during the second attempt were mainly due to page data loading issues that blocked access to operation widgets.
@@
Completion Time. Figure 15 shows that completing tasks with Prompt2Task takes significantly longer during the first use than when carried out manually (p-value<0.001). Figure 15 indicates that the completion time correlates positively with the number of steps involved during the first use of Prompt2Task, which is more pronounced than with manual methods and the second use of Prompt2Task. The time stems primarily from the system’s retrieval for online tutorials and exploration of operation sequences, which is tied to the efficiency limitations of the LLM. Notably, time decreases substantially on second use, as the system’s historical task repository eliminates the need for additional tutorial retrieval and exploration, thereby becoming significantly faster than manual methods (p-value<0.001).
@@
User Intervention. When using Prompt2Task, each new task required 0.92 user interventions (sd = 1.07) on average. Of these, 0.75 interventions per task (sd = 1.04) occurred during the operation mapping. Among the user intervention requests made by the Assessment Agent, 66.67% were deemed necessary. Notably, the second time a task was executed using Prompt2Task, user interventions were no longer required.
@@
Figure 16. Distribution of participant ratings for each statement in Table 15, assessing various aspects of Prompt2Task. Ratings range from 1 (strongly disagree) to 7 (strongly agree).
Subjective Feedback. Post-experiment interviews were conducted with participants, and their 7-point Likert scale responses are summarized in Figure 16. The participants’ ratings demonstrate that Prompt2Task was well-received across several dimensions, including low effort, ease of use, low time pressure, decision intelligence, interactive intelligence, interactive flexibility, satisfaction, and willingness to reuse. Additionally, the following insights were gathered from the interviews:
  • Preference for Forms of Intervention: Participants had different experiences with various forms of intervention and expressed distinct expectations for improvement.
    For the dialogue, most participants (11/12) were satisfied, finding it an effective way to express their intentions. P3 expressed surprise and satisfaction, “Once I taught it the process, next time I only needed to provide a few keywords to complete the task.” P1 found the context supplement very helpful, which allows for queries without the need for “careful consideration”. However, P1 also found the dialogues somewhat verbose, stating, “Occasionally, the agent would ask me for unnecessary parameter details. For example, asking me for the WeChat version number—I don’t even know it. I would prefer the agent to generate some reliable options for me to choose from.”
    All participants indicated that they preferred demonstration over instruction editing. P2 commented, “As long as the text isn’t too far off, I won’t modify it. Editing the text is less convenient than just demonstrating the operation later.”
    Regarding demonstrations, participants (10/12) commonly noted that most interventions were for locating small icons, which they accepted. For improvement, P1 suggested incorporating images into the step descriptions for clarity, similar to the online tutorials including screenshots. Some participants found that the agent’s prompts were sometimes incorrect, leaving them unsure about how to proceed. P8 mentioned, “The agent asked me to click ‘workshop’, but there was nothing like that on the page.” P4, an older user, remarked, “Adding steps is quite challenging for me; I hope the tutorials could be more accurate.”
    In addition to the above interaction methods, P3 expressed a desire for more intuitive communication methods, “Guess what I am trying to do and help me out when I am stuck in a certain step.”
  • Tolerance for Latency: Participants’ tolerance for latency varied depending on the individual, interaction state, and task. Most participants (10/12) acknowledged that some complex tasks understandably took longer. P5 noted, “Although the process took a bit of time, the real-time feedback made it feel like a learning experience, which was acceptably paced for me.” However, some participants expected lower latency for simpler tasks during their first use of Prompt2Task. P1 indicated that while he could accept the execution time, he hoped for faster response in chatting. P2 felt frustrated when the agent was slow in performing tasks she already knew how to complete. P4 remarked, “It felt like I had to wait a bit for each step, which made me feel like I was just idling, even though it was only a few seconds. Perhaps some animations or interactive elements could help improve the experience.” Fortunately, after using Prompt2Task a second time, all participants expressed satisfaction with the completion time.
@@
  • Evaluation of Intelligence and Expectations: All participants acknowledged the system’s intelligence. P3 specifically praised the feature that “aligns the smartphone’s operational page with each step description”. Similarly, P5 stated, “Before, I struggled with the tutorial’s first step not matching my current page. Now it is a significant improvement.”. Room for improvement was also noted. P9 advocated for a more “aggressive” execution approach, arguing the current system is slightly conservative. P11 noted that some steps for confirmation were unnecessary. P12 observed, “Tasks generally succeeded, but some took a roundabout path.”
@@
  • Application Scenarios: Many participants praised the system’s transparency and user-centered approach. P8 stated, “It gives me a feeling like I’m navigating through the tasks. It’s very interesting.” The utility of Prompt2Task in simplifying otherwise complex tasks was another strong point. P1 expressed, “I often ask my family how to use various functions on my phone, but it’s challenging to follow their instructions. Now I can just copy their responses and the system executes the steps for me.” Similarly, P12 shared, “I find new applications, especially those with complex functions, very daunting. It’s such a hassle to contact customer service and then follow their steps. Prompt2Task saves me from all that trouble.” P9 pointed out, “I can define a few routine tasks, like accounting after spending money. The agent can then automate these tasks for me.” P3 stated, “I often forget the steps, so the replay is very useful for me.”
@@ 7. Discussion
In this section, we delve deeper into the significance of knowledge and the applications of our proposed Prompt2Task. We also discuss the limitations of the system and potential future work.
@@ 7.1. Open and Accumulated Knowledge
Prompt2Task aims to convert information across modalities (from texts to operations), with its stages focusing on bridging gaps between them and minimizing information loss. In addition to the collaboration of its multi-agents, Prompt2Task’s effectiveness is derived from its open and accumulated knowledge, ensuring its broad applicability and efficiency.
Prompt2Task leverages massive online tutorials to address a broader range of user needs, surpassing existing work that relies on pre-collected datasets (Li et al., 2020a; Venkatesh et al., 2022; Vu et al., 2023). It supports functionalities deep within applications, which exceeds the capabilities of current assistants like Google Assistant and HUAWEI Celia. Moreover, Prompt2Task’s approach to mobile semantics is open-ended, enabling exploration beyond rigid adherence to given instructions. The open-ended approach allows Prompt2Task to handle diverse textual inputs and adapt to changing environments, significantly expanding the task automation’s application scope.
Prompt2Task’s knowledge base evolves through user usage and feedback. The evolution manifests in improved success rates, rising from 60.80% to 95.24%, and reduced completion time for repeated tasks (Figure 15). Notably, the accumulation of knowledge contributes to a reduction in user interventions. When similar task requirements arise again, Prompt2Task operates autonomously, only requiring user intervention for contextual parameter adjustments or significant interface updates. The accumulated knowledge empowers Prompt2Task to tackle increasingly complex and nuanced tasks, dynamically adapting to emerging user needs.
It is natural to consider how Prompt2Task handles outdated knowledge. While accumulated knowledge enhances Prompt2Task’s performance by improving efficiency (Cases 1 and 4 in Appendix B.3), reducing user interaction in future tasks (Cases 2 and 3 in Appendix B.3), and increasing success rates (Case 5 in Appendix B.3), it is not indispensable. When a significant app update causes the accumulated knowledge outdated, Prompt2Task’s multi-agent design can still work effectively. As shown in Figures 5, 7, and 8, outdated or missing knowledge only affects specific parts. Therefore, the agents still have the potential to make correct decisions. Upon successful execution, the accumulated knowledge is updated.
Table 13. “Outdated tutorials” is the result of using the 500 ground-truth step descriptions (100 step descriptions over 5 rounds of testing) as input. “Tutorials not found” is the evaluation of the entire pipeline, which used a total of 12,500 textual prompts (2,500 prompts over 5 rounds of testing) as input.
Outdated Tutorials Tutorials Not Found
Success Rate User Intervention Success Rate User Intervention
Prompt2Task 75/155 - 611/1448 -
Prompt2Task(AK) - - 467/967 -
Prompt2Task(minimal UI+AK) 140/155 0.75(sd = 1.08) 698/855 0.68(sd = 1.22)
Prompt2Task(UI+AK) 140/155 0.75(sd = 1.08) 91/110 0.86(sd = 1.21)
As illustrated in Table 13, Prompt2Task demonstrates high success rates even with outdated tutorials or when no tutorials are found, indicating that the Grounding Agent does not entirely rely on tutorials. In the worst-case scenario, where no tutorials match, no useful accumulated knowledge is available, and no human intervention is provided, Prompt2Task still achieves a 42.20% success rate. Even if the system fails, stage-wise transparency allows users to help the agent complete the task correctly.
@@ 7.2. Applications
By incorporating massive external knowledge and accepting various textual inputs, Prompt2Task supports a wide range of task executions, benefiting both end users and developers.
For end users, Prompt2Task improves the usability of smartphones and facilitates assistance between peers. It offers guidance for users to navigate unfamiliar smartphone functions, enabling them to access and automate tasks deep within applications. It allows users to personalize their smartphone experience by automating their own complex or repetitive tasks. Furthermore, textual assistance from various sources, such as family, friends, or online customer service, can be fed to Prompt2Task for automating tasks, which is valuable for teamwork or external support. For instance, an elderly user receiving textual assistance from family members can simply input the text into Prompt2Task for automating tasks, without the need to repeatedly switch between text and complex interfaces.
For developers, Prompt2Task can reduce the costs of developing and maintaining task automation. It can cost-effectively transform a vast array of existing tutorial texts into operations, seamlessly integrated into current voice assistants or online customer service, thereby expanding the range of supported tasks. Application and platform developers can leverage Prompt2Task to transform text-based cases in bulk into interactive tutorials or executable test scripts. It significantly reduces the cost of updating voice assistants, interactive tutorials, and test scripts, as developers need only update the source texts regularly to respond to the latest application functions and interfaces. Nonetheless, in sensitive applications such as banking, implementing robust authentication mechanisms is crucial to prevent malicious command execution.
Additionally, although Prompt2Task is designed for smartphones, its underlying methodology, particularly the multi-agent framework and extensive knowledge base, can be adapted to automate tasks in other domains, such as websites and IoT devices.
@@ 7.3. Limitations and Future Work
We outline the limitations of Prompt2Task, along with possible directions for future enhancements.
Adapting Dialogue Mechanisms for Diverse User Groups: The current design of the Analysis Agent offers a generalized approach without catering to the specific needs of different user demographics. For instance, older and younger users may express issues differently, necessitating a more nuanced, tailored dialogue system to accommodate their needs.
Parsing Complex Tutorial Logic: Automating responses to conditional branches like “If the version number is less than 14.0, please upgrade the app” could enhance the system’s efficiency. Future work will focus on more sophisticated logic analysis, including nested conditions.
Improving Semantic Understanding of GUI: In the present design, the mobile semantics module emphasizes the textual descriptions of mobile pages. However, current models often fall short, particularly when faced with pages not meeting accessibility standards. It results in blocking access to operation widgets and a notable loss of information during the translation from interface visuals to text. Moving forward, we aim to enhance algorithms that better recognize semantic structures within GUI, with the potential improvement of image recognition technologies.
Balancing Breadth with Depth: Currently, to expand the application scope, Prompt2Task applies the capabilities of LLM across various stages. However, some domains have not been deeply explored, such as addressing ambiguities in natural language (which the LLM’s rephrasing might not address) and setting instruction parsing templates to reduce the risk of LLM fabrication. As Prompt2Task’s usage data increases, we will systematically collect domain-specific data and user feedback to fine-tune the LLM or train local models, gradually reducing reliance on generalized templates and increasing its capacity for nuanced, domain-specific processing.